# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

### v4

@@ Abstract
Various industries such as finance, meteorology, and energy produce vast amounts of heterogeneous data every day. There is a natural demand for humans to manage, process, and display data efficiently. However, it necessitates labor-intensive efforts and a high level of expertise for these data-related tasks. Considering large language models (LLMs) showcase promising capabilities in semantic understanding and reasoning, we advocate that the deployment of LLMs could autonomously manage and process massive amounts of data while interacting and displaying in a human-friendly manner. Based on this, we propose Data-Copilot, an LLM-based system that connects numerous data sources on one end and caters to diverse human demands on the other end. Acting as an experienced expert, Data-Copilot autonomously transforms raw data into multi-form output that best matches the user’s intent. Specifically, it first designs multiple universal interfaces to satisfy diverse data-related requests, like querying, analysis, prediction, and visualization. In real-time response, it automatically deploys a concise workflow by invoking corresponding interfaces. The whole processes are fully controlled by Data-Copilot, without human assistance. We release Data-Copilot-1.0 using massive Chinese financial data, e.g., stocks, funds and news. Experiments indicate it achieves reliable performance with lower token consumption, showing promising application prospects.
Figure 1: Data-Copilot is an LLM-based system for data-related tasks, bridging billions of data and diverse user requests. It designs versatile interface tools for data management, invocation, processing, and visualization. Upon receiving a complex request, Data-Copilot autonomously invokes self-design interfaces and constructs a workflow to fulfill human intent. Lastly, raw data is converted into graphics, tables, and text forms tailored to human tastes.
@@ 1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. This wide-ranging, multiform data encapsulates critical insights that could be leveraged for a host of applications, from predicting financial trends to monitoring energy consumption.
Recently, the advancement of large language models (LLMs) (Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023), particularly the ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023a), has paved the way for advanced AI systems. Leveraging chain-of-thought (Wei et al., 2022c; Kojima et al., 2022; Gao et al., 2022b; Wang et al., 2023c), reinforcement learning from human feedback (Ouyang et al., 2022; Wang et al., 2022b), and instruction-following (Wei et al., 2022a; Iyer et al., 2022; Chung et al., 2022), LLMs demonstrate remarkable abilities in dialogue, reasoning and planning. However, in the face of the sheer magnitude and complexity of data, LLMs are confronted with the colossal challenge of managing, processing, and displaying data.
To achieve this, several challenges must be addressed: (1) From a data perspective: employing LLMs for directly reading and processing massive data is not only impractical but also poses the potential risks of data leakage. (2) From the model perspective: due to limitations of the input context, LLMs are unable to process large-scale data at once and are also not adept at handling numerical computations. (3) From the task perspective: many data-related tasks are intricate and require a combination of many operations, like retrieval, computations, and table manipulations, and present their results in multiple formats, e.g., images, tables, and text, which exceeds the capabilities of LLMs. These challenges significantly constrain the application of LLMs in data-related tasks.
Recently, many works have embarked on valuable explorations. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DIN-SQL (Pourreza & Rafiei, 2023) and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL. However, our focus is more on how to better process and present data based on human preference, rather than merely retrieving it from databases. Additionally, while SQL is convenient, it can not directly satisfy common data analysis needs such as prediction and visualization.
In light of this, we contemplate how to integrate LLMs into every facet of data-related tasks, not just as a Text2SQL solver. More specifically, we explore how LLMs can efficiently and autonomously process vast amounts of data to meet diverse human needs, without the necessity of direct data access. To pursue the answer, we trace back to the origins of data science in the 1970s. The pioneer of data science, Peter Naur (Turing Award winner in 2005), defined data science as follows:
Data science is the science of dealing with data and processing large amounts of data. Humans as sources and receivers of data. The data must be chosen with due regard to the transformation to be achieved and the data processing tools available. —Naur, 1974
His statement elucidates how humans accomplish data-related tasks: humans create various tools that are easy to use and then transform the data into a human-friendly format. Similarly, LLMs also should play a more intelligent role like humans: designing tools for data management, and invoking tools for processing and presenting data.
Additionally, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). These LLM-based agents showcase the potential of LLMs in meeting human needs through code as an intermediary.
Inspired by them, we advocate utilizing the coding capabilities of LLMs to create diverse and powerful tools, thereby facilitating the management and utilization of vast amounts of data. Based on this, we propose Data-Copilot, a reliable data agent that manages vast data and dynamically fulfills a wide spectrum of human needs. As shown in Figure 1, Data-Copilot autonomously designs versatile interface tools for data management, and dispatch these interface tools step by step for diverse human requests, e.g., data analyzing, forecasting, and visualization, forming a data-to-human workflow.
As an LLM-based agent, Data-Copilot is versatile. It seamlessly connects an array of data sources from various domains on one end, while adeptly catering to diverse user demands on the other. Furthermore, Data-Copilot is scalable. It can continuously extend its capability boundaries by developing new interface tools for emerging new data and requests. Our system comprises two processes:
  • Interface Design: Data-Copilot adopts an iterative self-request process to fully explore the data, mimicking a broader range of human need scenarios. As shown in Figure 1, starting from a few seed requests, it self-generates a large number of requests and then abstracts them into generalized interface tools for efficient utilization. These interfaces empower Data-Copilot to efficiently perform data querying, processing, forecasting, and visualization.
  • Interface Dispatch: When a user request is received, Data-Copilot first parses the user intention and then plans an interface invoking workflow for the user request. This step-by-step invocation can efficiently fulfill user requirements while minimizing token overhead.
Figure 2: Overview of Data-Copilot. Interface Design: We devise a self-request process allowing LLM to generate sufficient requests from a few seed requests autonomously. Then LLM iteratively designs and optimizes interfaces based on generated requests. These interfaces are described in natural language, making them easily scalable and transferable across different platforms. Interface Dispatch: Upon receiving user requests, LLM plans and invokes interface tools based on self-design interface descriptions and in-context demonstrations. This allows for the deployment of a logical workflow that fulfills user demands and presents the results to the user in multi-form.
Incorporating two phases, Data-Copilot achieves autonomous data management and utilization without the need for direct data access. Our contributions can be summarized as follows:
@@
    To efficiently handle data-related tasks, we design a universal system, Data-Copilot, that connects massive data sources to diverse user tastes. It leverages LLM for more efficient data management and utilization, reducing tedious labor.
@@
    Data-Copilot can autonomously manage, process, analyze, and visualize data. For a human request, it transforms raw data into informative results that best match the user’s intent.
@@
    As a reliable assistant, Data-Copilot abstracts a multitude of common needs into various practical interface tools. When faced with real-time requests, it invokes these related interfaces sequentially or in parallel for problem-solving.
  4. 4.
    We evaluate Data-Copilot on massive Chinese financial data, e.g., stocks, funds, macroeconomics, and news. The excellent success rate and lower computational cost showcase its potential and application prospects.
@@ 2 Data-Copilot
Data-Copilot is an LLM-based agent that accomplishes automated management, invocation, and processing of a large number of data by self-design sophisticated interface tools. Besides, Data-Copilot adeptly comprehends user intentions, and invokes well-designed tool interfaces for request solving, significantly reducing the need for tedious labor.
As shown in  Figure 2, the LLM plays the role of a designer (Section 2.1) to independently design various tool interfaces to for data management and utilization. Besides, LLM also acts as a dispatcher (Section 2.2), invoking generated interfaces and deploying the workflow for problem-solving. Through the integration of two stages, Data-Copilot manages to efficiently handle a large volume of data-related tasks and caters to various user needs.
2.1 Interface Design
Just as Peter Naur pointed out, designing a set of versatile interface tools that are easy to manage and use is critical. Therefore, Data-Copilot designs a plethora of interfaces as tools, where each interface is composed of natural language (functional description) and code (implementation), responsible for data acquisition, processing, and visualization.
As shown in Figure 2, beginning with some seed requests, LLM autonomously mimics and generates a large number of requests. Then LLM designs common interfaces for these requests, representing a set of universal needs. By gradually refining and expanding their functionalities, these interfaces progressively cover a more diverse array of needs. Ultimately, the LLM translates these interfaces into executable and callable code modules.
Explore Data by Self-Request
The design of the interface depends on two aspects: what kind of data is available and also what kind of demand the user proposes. Inspired by (Wang et al., 2022a; Dibia, 2023), Data-Copilot first autonomously explores the data to mine more requests. Then it proceeds to design the interfaces for these requests.
Specifically, we generate a parsing file for each data source to help LLM understand the data. Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example. This process does not require too much labor, as data providers usually offer a description of the data and access methods.
Then we adopt a self-request process to explore data, i.e. the parsing files and a few seed requests are fed into the LLM and then LLM is instructed to generate more diverse requests by mimicking. This process aims to explore an extensive array of data sources while covering diverse user requests. In Figure 2, LLM generates four new requests based on the two seed requests and data parsing files. The detailed prompts are shown in Section F.1.
Interface Definition
In this step, Data-Copilot creates various interface tools to fulfill the previously generated requests. Specifically, we feed all data parsing files and all interfaces stored in the interface library (empty at first iteration) into LLM as a prompt. LLM needs to resolve each request one by one. When addressing each request, LLM can use existing interfaces from the interface library or redefine a new interface to fulfill the task.
Concretely, if existing interfaces are insufficient to address the current request, LLM can create a new interface: Interface1={Name:"", Function description:"", Input and Output:""}... After that, LLM proposes a feasible solution using these self-design or existing interfaces: Solution="These three interface functions will meet your needs. You should first use the interface: {getBankData} to acquire data, then...". The prompts and cases are shown in Sections F.1 and G.1.
Importantly, Data-Copilot only uses natural language to define the interface’s functions and their arguments, without considering the specific implementation, which is grammar-free. This process is similar to software architecture design by a human programmer. It allows Data-Copilot to focus on designing the layout and relation of the interfaces with different functionalities, rather than complying with programming grammar.
Interface Merging
@@ 2.1 Interface Design
To make the interface more universal, Data-Copilot considers whether each newly designed interface can be merged with the existing ones in the library. Specifically, when a new interface is designed, Data-Copilot also checks whether this interface is similar to the previous ones, in terms of functionality, arguments, etc. Two similar interfaces are merged to create a new and generalized interface. As shown in Figure 2, two interfaces {Interface: Query-GDP} and {Interface: Query-CPI} are merged into one {Interface: Query-GDP-CPI}. This merging process is similar to software developers wrapping similar modules into a new one.
After this process, a large number of similar interfaces are merged. Each interface within the library possesses distinct functionalities, which are adeptly suited for real-time response applications.
Interface Implementation
Data-Copilot iteratively processes each request through interface definition and interface merging. Eventually, when all the requests can be satisfied by these self-design interfaces, Data-Copilot transforms these interface tools into executable and callable code modules.
The whole interface design process is offline. As shown in Figure G3, Data-Copilot automatically produces five types of interfaces in the interface library: data acquisition, processing, prediction, visualization, and DataFrame manipulation. It transforms repetitive and tedious labor into an automated process, and also can effortlessly add new interfaces for additional requests or new data sources. Besides, Data-Copilot easily switches to other programming platforms and databases by simply re-generating the implementation code for the interface, demonstrating excellent scalability.
2.2 Interface Dispatch
In the first stage, Data-Copilot designs a variety of generic interface tools for data acquisition, processing, and visualization. These interfaces are versatile and easy to use, improving the efficiency of real-world response. These interfaces are invoked in this phase and form an efficient workflow for real-time requests (Figure 2). A detailed prompt and cases are shown in Sections F.2 and G.2

### v5

@@ Abstract
Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily. Efficiently managing, processing, and displaying this data requires specialized expertise and is often tedious and repetitive. Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution. However, LLMs are not adept at handling complex numerical computations and table manipulations and are also constrained by a limited context budget. Based on this, we propose Data-Copilot, a data analysis agent that autonomously performs querying, processing, and visualization of massive data tailored to diverse human requests. The advancements are twofold: First, it is a code-centric agent that receives human requests and generates code as an intermediary to handle massive data, which is quite flexible for large-scale data processing tasks. Second, Data-Copilot involves a data exploration phase in advance, which explores how to design more universal and error-free interfaces for real-time response. Specifically, it actively explores data sources, discovers numerous common requests, and abstracts them into many universal interfaces (code modules) for daily invocation. When deployed in real-time requests, Data-Copilot only needs to invoke these pre-designed interfaces, transforming raw data into visualized outputs (e.g., charts, tables, and text) that best match the user’s intent. Compared to generating code from scratch, invoking these pre-designed and compiler-validated interfaces can significantly reduce errors during real-time requests. Additionally, interface workflows are more efficient and offer greater interpretability than code. We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news, demonstrating promising application prospects. Quantitative evaluations indicate that, compared to direct code generation and other agent strategies, our exploration-deployment strategy addresses human requests more accurately and efficiently, i.e., with higher success rates and lower token consumption.
Figure 1: We present two strategies for using LLMs to automate data analysis. Upper: Directly using LLMs to read and process massive data is limited by context length, the LLM’s capabilities, and output format constraints. Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks. It actively explores how to design more universal and error-free interface modules, improving the success rate of real-time requests. Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow to fulfill human intent.
@@ 1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption. However, these data-related tasks often require specialized knowledge and tedious manual labor.
Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022), especially when combined with techniques such (Wei et al., 2022b; Kojima et al., 2022) has demonstrated their capability to handle complex tasks. Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow that performs data analysis and outputs visualization results that best match user expectations?
An intuitive solution is to directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023), just as we do with text. However, several challenges must be considered: (1) Due to the input context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text. Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources. (2) Data processing is very complex, involving many tedious numerical calculations and intricate table manipulations. LLMs are not adept at performing these tasks. (3) An LLM can only output text, whereas we usually prefer charts or tables when analyzing data. These challenges significantly constrain the application of LLMs in data-related tasks.
Recently, many works have explored alternative solutions. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024) and DB-GPT (Xue et al., 2023) apply LLMs to data science domain like Text2SQL. Besides, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent. Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and output multi-modal results (e.g., chart, table, text) for humans, reducing tedious human labor and eliminating the need for direct data access by the LLM. However, creating a reliable code agent for various data analysis tasks is also not an easy feat. Firstly, LLM-generated code often contains formatting errors, logical inconsistencies, and code hallucinations, such as fabricating non-existent functions, which may prevent it from being deployed in real-world environments. Secondly, most code-based agent generates code for solving user requests, without considering the relations between different requests and only addressing each issue independently.
Based on this observation, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis. First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches the user’s intent. The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks. Besides, Data-Copilot also incorporates a data exploration phase. During this phase, it actively explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance. In a real-world deployment, Data-Copilot flexibly invokes pre-design interface modules to handle user requests, creating an interface workflow that transforms raw data into desired visualization results.
This design can bring three advantages. Firstly, this exploratory process allows Data-Copilot to analyze connections between data, design general interfaces for similar requests, and pre-validate their correctness, thereby reducing errors in real-time responses. Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving generation efficiency. Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
To achieve this, as shown in Figure 1, Data-Copilot contains three steps for data exploration:
@@
    Request Exploration: Data-Copilot explores potential requests and mines a broader range of human need scenarios from data. It involves a "self-instruct" process to generate massive simulated requests based on all data schemas (types and formats) and seed requests.
@@
    Interface Design: Our agent generates code for each request and then considers whether these requests have similar demands and whether general interfaces can be abstracted.
@@
    Interface Optimization: To improve the quality of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces by reflecting on feedback from the compiler.
After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different request scenarios, Data-Copilot can flexibly design different interface invocation procedures, such as serial workflows, i.e., step-by-step invocation, parallel workflows, and loop workflows. Our contributions are threefold:
Figure 2: Overview of Data-Copilot. Data Exploration: It devises a self-request process to generate numerous potential requests based on data schemas. Then it interleaves to design and optimize the interfaces to solve these requests. The whole exploration-design-optimization process is operated in advance, generating many general and error-free interface modules, such as data retrieval, computation, and visualization. Workflow Deployment: Upon receiving real-time requests, LLM invokes these interfaces, forming an interface workflow with different structures for request solving.
  1. 1.
    We propose a code-centric agent, Data-Copilot, for massive data analysis tasks. It leverages LLM and its code generation capabilities to automate data querying, processing, and visualization tailored to human requests, reducing tedious human labor.
  2. 2.
    We decouple code generation into two phases: data exploration and workflow deployment. In the exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data. These interfaces enhance the success rate and efficiency of real-time responses. In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests.
  3. 3.
    We open-source Data-Copilot for diverse Chinese financial data analysis, including stocks, funds, and live news. Quantitative evaluations indicate that our agent outperforms other agent strategies, with higher success rates and lower token consumption. Besides, interface workflows are more convenient for human inspection and interaction, offering stronger interpretability.
@@ 2 Data-Copilot
Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent. Data-Copilot operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-request process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these "self-instruct" requests, it generates code and abstracts universal interface modules for each request (Step 2: design). After that, it optimizes similar interfaces and reflects on the code compiler feedback, ensuring each interface is universal and correct (Step 3: optimization). Finally, Data-Copilot acquires many generic, error-free interfaces for subsequent use. The whole exploration-design-optimization process is operated in advance.
When faced with real-time requests (Section 2.2), Data-Copilot adeptly comprehends user intentions, and invokes pre-designed interfaces for request solving, significantly reducing the need for tedious labor. We provide a detailed prompt for two phases in Section F.1 and F.2.
2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe what data is available, understand the data formats, and determine how to use them. Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability. Similar to humans, Data-Copilot aims to explore data and derive insights from vast data sources. These insights include understanding the relations between the data, discovering the potential requests associated with the data, and organizing the code framework efficiently. Based on this, Data-Copilot summarizes these exploratory insights into many reusable code components and optimizes its performance. It significantly reduces the code complexity for real-time responses, thus improving the success rate.
Request Exploration
To explore data and mine insights, we design a self-request process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs must generate requests specifically related to these data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.
To achieve this, we first generate a parsing file for each data source to help LLM understand the data. Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example. Then we feed the parsing files and a few seed requests into the LLMs, and prompt LLMs to generate more diverse requests based on these data, covering different demand scenarios as comprehensively as possible. The brief prompt is shown as follows:
###Seed Request: {request1, request2,...}
###Parsing file for GDP_Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: pro.cn-gdp(start-time, end-time, frequency,...),
Output Schema: Return 9 columns, including quarter, gdp, gdp-yoy...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’,..),
First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
###Parsing file for Stock_data:{...},
How to ensure the quality of the generated requests? A common issue is that the LLM often proposes a request about the data that doesn’t exist. To address this, we design a backward verification process to check the rationality of these requests. Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests. Upon manually check some requests, we find that the generated requests generally met our expectations.
Interface Design
After generating massive requests through the self-request, Data-Copilot considers abstracting universal interfaces from these requests. First, we need to define what an interface is in our paper. Similar to human programming practices, we refer to a function module abstracted from the code as an interface. Each interface is composed of a name, parameters, a functional description, and the implementation code, performing specific functions such as data retrieval, computing, and plotting. To abstract general interfaces from all requests, we design an interwoven process, performing interface design and interface optimization alternately for each request.
Firstly, we feed related data parsing files and existing interfaces stored in the interface library (empty at first iteration) into LLMs as a prompt. LLMs need to generate a complete code for each request. When generating, LLMs are instructed to prioritize using the existing interfaces from the previous iterations. If these interfaces are insufficient to address the request, the LLMs design a new interface. As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...). As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
Interface Optimization
To improve the quality of interfaces, Data-Copilot optimizes them in two ways: merging similar interfaces to make them more universal and reflecting on compiler feedback to reduce errors. This optimization process is intertwined with the interface design for each request.
@@ 2.1 Data Exploration
Firstly, similar to software developers wrapping similar code modules to reduce redundancy, Data-Copilot considers whether each newly designed interface can be merged with the previous interfaces. LLMs are instructed to check the functionality, parameters, and logic of the new and existing interfaces, and merge those similar interface modules to create a new one. As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...). This interface optimization process makes each interface more general and rational.
Besides, to reduce code hallucinations or formatting issues, Data-Copilot also self-reflects on the feedback of code execution for each request if compilation errors occur. We have LLMs perform self-reflection until the code passes or the maximum retry number is reached. This process ensures the correctness and usability of each interface.
For each request, Data-Copilot alternates the above two steps: interface design and optimization until all the requests can be satisfied by these interfaces. Finally, as shown in Figure G3, Data-Copilot produces five types of interfaces: data acquisition, processing, prediction, visualization, and DataFrame manipulation. We provide some detailed cases in Section G.1. Besides, to handle newly added data sources, Data-Copilot can repeat the aforementioned exploration-design-optimization process on new data. It can create some new interface modules for these new data.
2.2 Workflow Deployment
In the first phase, Data-Copilot designs many generic interfaces with different functionalities. They are versatile and easy to use, simplifying the code generation of real-time responses. As shown in Figure 2, when faced with human requests, Data-Copilot only needs to invoke related interfaces and form an interface workflow. A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex call flow diagram. This workflow transforms raw data into visualized results. The detailed prompts and cases are shown in Sections F.2 and G.2.